OpenStudy (anonymous):

HELP: Solve by Factoring. 9x^2+4x=0

OpenStudy (anonymous):

9x^2+4x=0 x(9x+4)=0 either x=0 or, 9x+4=0

OpenStudy (anonymous):

would it be (3x+2)(3x-2) do that is 6x-6x=0?

OpenStudy (anonymous):

I'm not sure your answer is correct

OpenStudy (anonymous):

we can not factorize it like this because a^2-b^2=(a+b)(a-b)

OpenStudy (anonymous):

if the question is like 9x^2-4 then we do this

OpenStudy (anonymous):

but we would multiply the inner product and then the outer product and then add together. In my example ... I know that is one of the ways on my homework. I was just unsure.

OpenStudy (anonymous):

in the question we see that there is "x" common so, we first take out "x" 9x^2+4x = x(9x+4)

OpenStudy (anonymous):

so, the factorize form of 9x^2+4x is x(9x+4)

OpenStudy (anonymous):

yes that's right sorry read it wrong

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

now either x=0 or , 9x+4=0 so, x=0 is one of the answer and we solve 9x+4 =0 for other answer

OpenStudy (campbell_st):

the only common factor is x so you have x(9x + 4) = 0 then x = 0 or -4/9 as the solutions to the equations